# Revisiting the relaxation of constraints in gauge theories

### v3

@@ 1 Introduction
Of course, we should try all possible theoretical ideas, and the constraints’ relaxation is no exception to that. However, these papers claimed it to be a necessity [1, 2] of quantum physics based on a rather naïve treatment of path integrals as if, modulo summing over the gauge copies, those were always well-defined and mathematically clean which would probably entail a perfect quantum gravity theory with flawless background independence and everything, let alone the mathematical rigour of all our quantum field theories. This is not the case.
Another thing to mention is that, with the poor level of rigour typical of physicists, we do know how to quantise gauge theories without relaxing the constraints, and even in a few different approaches. We recommend a very nice paper [5] and references therein for a review on the topic. Note that another point they make [5] is also worth paying attention to. Namely, these ideas of relaxing the constraints, though admitting that it is a modification at the level of the classical theory, have been appearing many times in the literature and date back to old works by V.A. Fock and E.C.G. Stückelberg; see the review and references in the paper [5].
@@ 4 The general arguments for mechanical systems
An important point about electrodynamics above is that its gauge freedom ”hits twice” [6]. The gauge transformation mixes different types of derivatives. The $A_{0}$ component lacks the time derivative in the action, while another component, $A_{i}$, lacks it own $\partial_{i}$ derivative there. Therefore, the primary constraint is not a gauge generator in itself, unless we go for the extended Hamiltonian, and it generates the secondary constraint which can finally be translated into the Lagrangian one. Only their combination then serves as a generator of gauge transformations in the initial meaning [14].
@@ 5 The case of gravity
Finally, let us make some comments about relaxation of constraints in vacuum general relativity. In this case, there are immediately two different choices of variables for Hamiltonian description, the metric tensor components and ADM variables [16]. One way is to use the metric tensor components à la Dirac [17], with the primary constraints of $\pi^{0\mu}=0$ after some transformation of the action functional. Relaxation gauge conditions of $g_{0\mu}=f_{\mu}(x)$ make the action, or the total Hamiltonian, produce then [2] only the spatial part of Einstein equations $G^{ij}=0$. Therefore, it is indeed akin to adding an effective energy-momentum tensor with only non-spatial components $T^{0\mu}$ non-vanishing [2]. Of course, it is subject to covariant conservation.
@@
However, in general for fixing non-zero shifts, it is not the case. Even the Hamiltonian analysis itself gets some modification due to changing the variables to ADM. It makes some people claim that those variables are not good for a proper Hamiltonian analysis of general relativistic dynamics [21, 22]. In reality, there is no problem, and it is just a feature of any gauge theory where even the canonical Hamiltonian is not uniquely defined [8].

### v4

@@ 1 Introduction
Of course, we should try all possible theoretical ideas, and the constraints’ relaxation is no exception to that. However, these papers claimed it to be a necessity [1, 2] of quantum physics based on a naïve approach to path integrals as if, modulo summing over the gauge copies, those were always well-defined and mathematically clean which would probably entail a perfect quantum gravity theory with flawless background independence and everything, let alone a consistent mathematical description of all our quantum field theories, nonperturbatively. This is not the case.
Another thing to mention is that, with the poor level of mathematical rigour available in quantum field theory, we do know how to quantise gauge theories without relaxing the constraints, and even in a few different approaches. We recommend a very nice paper [5] and references therein for a review on the topic. Note that another point they make [5] is also worth paying attention to. Namely, these ideas of relaxing the constraints, though admitting that it is a modification at the level of the classical theory, have been appearing many times in the literature and date back to old works by V.A. Fock and E.C.G. Stückelberg; see the review and references in the paper [5].
@@ 4 The general arguments for mechanical systems
An important point about electrodynamics above is that its gauge freedom ”hits twice” [6]. This expression means that one gets twice as many first-class constraints as there are independent gauge symmetries, which in turn leads to elimination of two dynamical modes for every such symmetry. Therefore, with only one gauge symmetry of electrodynamics, there are 42=24-2=2 dynamical modes. The gauge transformation mixes different types of derivatives. The $A_{0}$ component lacks the time derivative in the action, while another component, $A_{i}$, lacks it own $\partial_{i}$ derivative there. Therefore, the primary constraint is not a gauge generator in itself, unless we go for the extended Hamiltonian, and it generates the secondary constraint which can finally be translated into the Lagrangian one. Only their combination then serves as a generator of gauge transformations in the initial meaning [14].
@@ 5 The case of gravity
Finally, we would like to comment on relaxation of constraints in vacuum general relativity (GR). Before we do so, please note that one might say that general relativity is not a usual gauge theory, and therefore it is irrelevant for the gauge theory discussions. It is then possible to turn to other formulations such as MacDowell-Mansouri [16] or new works on Lorentz gauge theory [17, 18, 19], or to mathematical considerations of teleparallel gravity [20, 21]. However, for our purposes, GR is a gauge theory, for it does enjoy a local symmetry when formally taken in terms of the metric tensor components as its variables.
In this case, there are immediately two different choices of variables for Hamiltonian description, the metric tensor components and ADM variables [22]. One way is to use the metric tensor components à la Dirac [23], with the primary constraints of $\pi^{0\mu}=0$ after some transformation of the action functional. Relaxation gauge conditions of $g_{0\mu}=f_{\mu}(x)$ make the action, or the total Hamiltonian, produce then [2] only the spatial part of Einstein equations $G^{ij}=0$. Therefore, it is indeed akin to adding an effective energy-momentum tensor with only non-spatial components $T^{0\mu}$ non-vanishing [2]. Of course, it is subject to covariant conservation.
@@
However, in general for fixing non-zero shifts, it is not the case. Indeed, in the purely metric component variables, the standard relaxation goes in terms of restricting the variations to $\delta g_{00}=\delta g_{0i}=0$, thus loosing the non-spatial part of Einstein equations. At the same time, the ADM-type relaxation would make the lapse and shift fixed, and that means $\delta g_{0i}=0$ and $\delta g_{00}=-N_{i}N_{j}\delta h^{ij}$ such that, in case of fixing a non-vanishing shift, we get a nontrivial combination of spatial equations with the temporal one, similar to what we have seen above when fixing a gauge of $A_{0}+f(A_{i})=0$ in the Maxwell action.
Even the Hamiltonian analysis itself gets some modification due to changing the variables to ADM. It makes some people claim that those variables are not good for a proper Hamiltonian analysis of general relativistic dynamics [27, 28]. In reality, there is no problem, and it is just a feature of any gauge theory where even the canonical Hamiltonian is not uniquely defined [8].